Shack Homebuilders Limited is evaluating a new promotional campaign that could increase home sales. Possible outcomes and probabilities of the outcomes are shown next. Additional Sales in Units 40 90 115 Possible Outcomes Ineffective campaigrn Normal response Extremely effective Probabilities 30 .50 20 Compute the coefficient of variation. (Do not round intermediate calculations. Round your answer to 3 decimal places.) Coefficient of variation
Answer #1

Expected Sales=Respective Sales*Respective probability

=(0.3*40)+(0.5*90)+(0.2*115)=80 units

probability Sales probability*(Sales-Expected Sales)^2
0.3 40 0.3*(40-80)^2=480
0.5 90 0.5*(90-80)^2=50
0.2 115 0.2*(115-80)^2=245
Total=775

Standard deviation=[Total probability*(Sales-Expected Sales)^2/Total probability]^(1/2)

=27.84 units(Approx)

Coefficient of variation=Standard deviation/Expected Sales

=(27.84/80)=0.348(Approx).